OpenStudy (anonymous):

Our class planned a holiday party for disadvantaged kids.some of us baked cookies for the party .on the day of the party we found we could divide the cookies into packets of two,three,four,five,or six and have just one cookie leftover in each case.if we divided them into packets of seven there would be no cookies left over.what is the least amount of cookies the class could have baked?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

It's some multiple of 7 for sure.

OpenStudy (anonymous):

Yea and the number is odd too

OpenStudy (phi):

what are you studying?

OpenStudy (anonymous):

This Is algebra

OpenStudy (anonymous):

R u there

OpenStudy (phi):

yes, but so far nothing is coming to mind.

OpenStudy (phi):

if it's multiple choice, we could test each choice. or if we *really* needed an answer, we can have a computer search for an answer. But I think they want a clever way (which you have been studying??) what have they been teaching that relates to this question?

OpenStudy (anonymous):

We'll today we did a warm up on Order of operations

OpenStudy (phi):

let me think about it.

OpenStudy (anonymous):

Ok

OpenStudy (phi):

301= 3*43 works but I don't know a quick way to find it. Checking is easy: 301/2 = 150 r 1 301/3= 100 r 1 301/4 = 75 r 1 301/5 = 60 r 1 301/6= 50 r 1 301/7 = 43 r 0

OpenStudy (anonymous):

Wait why use 3*43

OpenStudy (anonymous):

@phi

OpenStudy (phi):

oh, I meant 7*43 = 301

OpenStudy (anonymous):

Where did you get 43 from

OpenStudy (anonymous):

So you divided 7 to 43

OpenStudy (phi):

But there must be a nicer way.

OpenStudy (anonymous):

Oh wait what would I do next

OpenStudy (phi):

what I tested were the multiples of 7: 0 7 14 21 28 35 42 49 56 63 70 77 84 91 98 105 and so on... when I got to 301, it worked.

OpenStudy (anonymous):

And to find out it worked you divided 301 to 7 and got 43 which is the least number of cookies

OpenStudy (phi):

I should not have mentioned the 43 (that is the number of packets with 7 cookies in a packet) the answer is 301 is the minimum number of cookies that they baked

OpenStudy (anonymous):

Oh ok thanks so all I had to do was find the multiple of seven and since the other numbers had variables they didn't work so 301 is my final answr

OpenStudy (phi):

It was a lot of work to do: check each multiple of 7 gives a remainder of 1 when divided by 2,3,4,5,6 I checked about 43 numbers (we could have eliminated the even numbers, so only checking about 20 something numbers.... but still a lot of work...

OpenStudy (anonymous):

Oh thank you